OpenStudy (anonymous):

Where is the first derivative of y = 3xe−x equal to 0?

OpenStudy (anonymous):

okay so iv got the derivative (-3e^(-x))(x-1)

jimthompson5910 (jim_thompson5910):

So they gave you \[\Large y = 3xe^{-x}\] right?

OpenStudy (anonymous):

yes

jimthompson5910 (jim_thompson5910):

I'm getting \[\Large \frac{dy}{dx} = -3e^{-x}(x-1)\] as well

jimthompson5910 (jim_thompson5910):

now set dy/dx equal to 0 and solve for x

OpenStudy (anonymous):

yea well thats what i cant do

jimthompson5910 (jim_thompson5910):

\[\Large \frac{dy}{dx} = -3e^{-x}(x-1)\] \[\Large 0 = -3e^{-x}(x-1)\] \[\Large -3e^{-x} = 0 \text{ or } x-1 = 0\]

jimthompson5910 (jim_thompson5910):

I'm using the idea A*B = 0 means either A = 0 or B = 0

geerky42 (geerky42):

Does that make sense? @Jdosio

OpenStudy (anonymous):

okay sorry just fell asleep in my desk yes iv got it thank you verry much:D